OpenStudy (togundeji):

19. Calculate the hydrogen-ion concentration [H+] for the aqueous solution in which [OH–] is 1 x 10–11 mol/L. Is this solution acidic, basic, or neutral? Show your work.(3 points)

563blackghost (563blackghost):

[H+] and [OH-] = 1 x 10^-14 Since we know that [OH-] is 1 x 10^-11 we would divide... \(\Huge{\frac {[OH-] [ H+] }{[OH-]}}\)

563blackghost (563blackghost):

So... \(\Huge{\frac{1 \times 10^{-14}}{1 \times 10^{-11}}=[H+]}\)

563blackghost (563blackghost):

If [H+] is greater than [OH-] then it would be acidic if it is not then it would not be acidic...

563blackghost (563blackghost):

Did you understand as to what I typed?

OpenStudy (togundeji):

1 x 10^-3 = 3 < 7 pH=3 <7 solution is acidic. I am wondering if this is right? @563blackghost

563blackghost (563blackghost):

Correct :) The [H+] is 1 x 10^-3 as well as acidic ^^

OpenStudy (togundeji):

Thank you! @563blackghost

563blackghost (563blackghost):

np :)